OpenStudy (anonymous):

Given a rectangle whose perimeter is 48 inches. The length of the rectangle is 4 inches longer than the width. What are the length and width of this rectangle. (The formula for the perimeter of a rectangle is P = 2l + 2w)

OpenStudy (anonymous):

Length = 14 inches and Breadth/Width = 10 inches.

OpenStudy (anonymous):

Length = l Breadth = b 2(l+b) = 48 l+b = 24----> Equation 1 Equation 2: l = b + 4 Now, substitution Equation 1 in Equation 2 : b + 4 + b = 24 2b = 20 Therefore, b = 10. And substituting the answer in Eq.2, we get l= 10 + 4 which is 14. Therefore, l = 14 and b = 10.